Adult tickets for a play cost $18 and child tickets cost $15. If there were 27 people at a performance and the theater collected $429 from ticket sales, how many children attended the play?
Answer by checkley71(8403)

18a+15c=429
a+c=27 or c=27-a
now replace c in the first equation with (27-a) & solve for a.
18a+15(27-a)=429
18a+405-15a=429
3a=429-405
3a=24
a=24/3
a=8 adult tickets sold.
27-8=19 children tickets sold.
proof:
18*8+19*15=429
144+285=429
429=429